OpenStudy (anonymous):

What is the ratio of the radius of a circle to its circumference? A. 1/2π B. 1/πr C. 1/π D. 2/π

OpenStudy (science0229):

\[2\pi r=c\]

OpenStudy (science0229):

right?

OpenStudy (science0229):

r=radius c=circumference

OpenStudy (anonymous):

wait which one would that be though?

OpenStudy (science0229):

I'll derive in 10 seconds.

OpenStudy (anonymous):

You know that pi is the ratio of the circumference to radius so radius to circumference is the inverse

OpenStudy (science0229):

Ratio of radius to circumference is\[\frac{ r }{ c }=\frac{ r }{ 2\pi r }=\frac{ 1 }{ ? }\]

OpenStudy (anonymous):

\[\frac{ radius }{ circumference }=\frac{ 1 }{ \pi }\]

OpenStudy (science0229):

correction:\[\frac{ 1 }{ 2\pi }\]

OpenStudy (anonymous):

2pi?

OpenStudy (anonymous):

thank you guys

OpenStudy (anonymous):

your awesome

OpenStudy (science0229):

FYI pi is the ratio of Diameter to circumference, not radius.

OpenStudy (science0229):

So 1/2pi is the correct answer.

OpenStudy (anonymous):

@sugahead12 @science0229 is right. I eant diameter not radius so it is 1/(2pi)